Be Careful Group has organized a social distancing gathering on September 16, 2020.

The cost estimates for the event are listed below.

Expected attendees: 40

Food: $1,200

Decorations: $100

Labor: $600

Hall Rental: $200

The actual costs are listed below.

Actual attendees: 50

Food: $1,450

Decorations: $105

Labor: $800

Hall Rental: $200

(1) Prepare a performance report with variances.

(2) Which items would you choose to investigate?
Answer #1

Statement of variance

Particulars Estimated cost Actual cost Standard cost for actual input Variance(Std cost-actual cost)
Food(variable) 1200 1450 (1200/40)*50=1500 50 favorable
Decoration(variable) 100 105 (105/40)*50=131.25 26.25
Labor 600 800 (600/40)*50=750 50 unfavarable
Hall rental(fixed) 200 200 200 0